# Improving probabilistic forecasts of extreme wind speeds by training statistical post-processing models with weighted scoring rules

### v2

@@ 6 Forecast combinations
These combine training using a CRPS and twCRPS and correspond to a weighted CRPS with weighting function $w(z)=1+\gamma\cdot\mathbb{1}(z\geq\tau)$, originally proposed by Thorarinsdottir and Schuhen (2018) for the assessment of forecasts. Other, non-threshold-based, weighting functions might also be usable to balance body-tail forecast performance, however indicator weighting has the advantage of closed-form solutions being available. In a second step we consider linear pooling between the CRPS-trained and twCRPS-trained model, to control twCRPS improvement and CRPS loss. We focus on the 80th percentile threshold for both truncated normal and truncated logistic distributions.
@@ 6.1 Weighted combinations
We consider training using weighted scores of the form given in eq. 16. In Figure 6 we show the CRPS skill over the CRPS-trained model (blue) and twCRPS skill over the twCRPS-trained model (orange) for different values of $\gamma$. For $\gamma=0$ the model corresponds to the CRPS-trained model whilst for growing $\gamma$ the twCRPS has stronger influence.
Varying $\gamma$ seems to allow users to control the CRPS loss and twCRPS improvement, with as $\gamma$ grows the resulting model approaching the twCRPS improvement of the twCRPS-trained model and its CRPS loss. Interestingly however, even though the twCRPS for the model with $\gamma=20$ is nearly at the level of the twCRPSt-trained model (at least for the truncated normal distribution), the CRPS skill loss is still much better with only -1.94% CRPS skill loss for the $\gamma=20$ model against -6.35% for the twCRPS-trained models in the truncated normal case and only -2.03% against -11.47% in the truncated logistic case (see section 5.1). Figure 15 in the appendix shows the twCRPS and CRPS for selected values of $\gamma=2,10,20$ for the truncated normal EMOS for both 80th and 90th thresholds. Combining CRPS and twCRPS through weighted training seems to avoid many "outlying" cases where CRPS performance falls off drastically through twCRPS training, whilst still capturing much of the twCRPS improvement. This is similar for the truncated logistic distribution (not shown).
When evaluating the performance of weighted training over a range of evaluation thresholds (figure 16(a) in the appendix for the truncated normal predictive distribution) one finds that the $\gamma$ parameter essentially controls the slope of the curve of twCRPS skill (over the CRPS-trained model): higher values of $\gamma$ lead to forecasts which are worse for the body of the distribution (measured in terms of twCRPS skill for low evaluation quantiles), but better in terms of twCRPS for higher evaluation quantiles, whilst small values of $\gamma$ control the twCRPS loss for small evaluation quantiles much more tightly, but also do not show as substantial of an improvement for high evaluation quantiles.
The parameter $\gamma$ presents a "hyperparameter" that can be chosen by users to tailor a model’s CRPS and twCRPS performance for specific applications. However after training, changing the balance of twCRPS improvement and CRPS loss requires retraining of the post-processing model. Thus in the following we will consider direct combination of CRPS-trained and twCRPS-trained models.
Figure 6: Weighted training: twCRPS skill score over the twCRPS-trained model (orange) and CRPS skill score over the CRPS-trained model (blue) as a function of $\gamma$ for both truncated normal (left) and truncated logistic (right) predictive distributions. The bands represent 95% intervals over stations.
@@ 6.2 Linear pool
The $\gamma$ parameter seems to "interpolate" between the CRPS-trained and twCRPS-trained model. We now study combining these models directly using their predictive CDFs. More concretely we study the following convex combination of the predictive distribution functions:
$F_{\text{pool}}=\lambda F_{\text{CRPS}}+(1-\lambda)F_{\text{twCRPS}}\rightarrow Z_{\text{pool}}\sim\begin{cases}Z_{\text{CRPS}}&\text{ if }U\leq\lambda\\ Z_{\text{twCRPS}}&\text{ if }U>\lambda\end{cases}$ (17)
@@
where $U\sim\text{Unif}(0,1)$ and $Z_{\text{CRPS}},Z_{\text{twCRPS}}$ follow $F_{\text{CRPS}}$ and $F_{\text{twCRPS}}$ respectively. This is also known as linear pool (Gneiting and Ranjan, 2013; Baran and Lerch, 2018) and the canonical case $\lambda=1/2$ corresponds to simple CDF averaging. The linear pool has some drawbacks, for example Gneiting and Ranjan (2013) show that it necessarily increases dispersion. More advanced versions such as the spread-adjusted linear pool have been proposed, however we defer their study for future work.
@@
We study the twCRPS skill over the twCRPS-trained model and the CRPS skill over the CRPS-trained model, whilst varying $\lambda$ in Figure 7. The $\lambda$ parameter allows to strike a balance between the CRPS (being best for the CRPS-trained model and worst for the twCRPS-trained model) and the twCRPS (vice-versa). The curves are somewhat similar to the ones obtained by weighted training. As one can see the twCRPS improvements and CRPS losses have however very different "steepness", meaning that for example a model with $\lambda=0.6$ still has most of the twCRPS improvements, but controls the majority of the CRPS loss. Evaluating the linear pool over a range of different evaluation quantiles (see figure 16(b) in the appendix for the truncated normal predictive distribution) we find, similar to the weighted training, that the $\lambda$-parameter allows to control the slope of the twCRPS skill curve (over the CRPS-trained model), meaning that for small values of $\lambda$ the twCRPS skill improvement in high quantiles and the twCRPS skill loss in low quantiles is larger, whilst both are smaller for values of $\lambda$ near $1.0$. For different applications different choices of $\lambda$ will be relevant.
@@
Figure 7: Linear pool: twCRPS skill score over the twCRPS-trained model (orange) and CRPS skill score over the CRPS-trained model (blue) as a function of $\lambda$ for both truncated normal (left) and truncated logistic (right) distributions. The bands represent 95% intervals over stations.
@@
We investigate the example case $\lambda=0.6$ for the truncated normal distribution. Figure 8 shows the twCRPS and CRPS skill (relative to the CRPS-trained model) for both the twCRPS-trained and the linear pool model. As one can see the linear pool model keeps much of the twCRPS improvement, but controls the CRPS loss much stronger, with many outlying cases not present. Pooling CRPS-trained and twCRPS-trained models seems to be able to somewhat alleviate the CRPS-twCRPS trade-off.
@@
Figure 8: Linear pool: twCRPS skill score (left) and CRPS skill score (right), both relative to the CRPS-trained model, for the twCRPS and the linear pool model with $\lambda=0.6$. Truncated normal predictive distribution, 80th percentile threshold.
@@
It might be possible to choose the $\lambda$ parameter in a data driven way, possibly based on the ensemble mean or external covariates such as weather patterns (Allen et al., 2021a; Spuler et al., 2024a). Simple experiments parametrizing $\lambda$ in terms of the distance of the ensemble mean to the threshold (a possible measure of the probability of exceedance), however did not show substantial skill over the ad-hoc choice of for example $\lambda=0.6$.
@@
Compared to the weighted training the linear pool directly interpolates between the two models, which means that for low values of $\lambda$ it is able to capture more of the twCRPS improvement, but also has much stronger CRPS loss. The latter is maximal with -1.9% for $\gamma=20$ for the weighted training, whilst reaching -19.5% for $\lambda=0$ in the linear pool case. By choosing adequate values of $\lambda$ and $\gamma$ however it is possible to construct models with very similar twCRPS and CRPS performance using both the linear pool and weighted training. Which one is preferable for applications might therefore depend on individual forecaster’s considerations, most notably whether changing the twCRPS/CRPS balance might be required during application. This would need retraining in the case of weighted training, making the linear pool preferable. For choosing $\gamma$ or $\lambda$ in applications there might be requirements on CRPS/twCRPS performance necessitating specific choices. Otherwise we recommend an ad-hoc choice of a value that somewhat balances the objectives.
@@ 7 Interpreting the training impact of the twCRPS
After experiments studying the training of post-processing models using the threshold-weighted CRPS we now consider a number of synthetic experiments to understand the training impact of the twCRPS better in an idealized system. We focus on deriving tail properties from simulated data and on extreme value distributions, which also have been used for various post-processing applications (see for example Baran and Lerch, 2015; Thorarinsdottir et al., 2016), to highlight how using a twCRPS can focus parameter inference on parameters controlling the distribution tail.
@@ 7.1 Background: Tail structures and proper scoring rules
Taillardat et al. (2023) argue that the mean (weighted) CRPS cannot distinguish between non-tail-equivalent forecasts. They show that for any random variable $Y$ and any $\nu>0$ it is always possible to construct a non-tail-equivalent random variable $X$ (meaning a random variable with different tail index and thus differing tail behavior/return levels for extremes), such that:
$|\mathbb{E}_{y\sim Y}\text{wCRPS}(F_{y},y)-\mathbb{E}_{y\sim Y}\text{wCRPS}(F_{x},y)|\leq\nu$ (18)
where wCRPS is the weighted CRPS with any weighting function $w$. Brehmer and Strokorb (2019) generalise this and show that this is not only a property of the weighted CRPS, but rather all proper scoring rules fail to distinguish between tail equivalent forecasts in the sense of Taillardat et al. (2023). As Brehmer and Strokorb (2019) conclude this casts doubt on the ability of proper scoring rules to distinguish between tail regimes. In practical applications however proper scores such as the log-score (negative log-likelihood) are often successfully used to fit distributions to extremes and to distinguish between tail structures. This is supported by the theory of M-estimation, of which scoring rule inference is a special case (Dawid, 2007), providing consistency guarantees on parameter estimates. Thus the authors of this work believe that the arguments presented in Taillardat et al. (2023) and Brehmer and Strokorb (2019) do not carry conclusive practical implications. This is especially as Taillardat et al. (2023) do not provide bounds on the constructed non-tail-equivalent random variables $X$ that show that these will influence statistical analysis in practice  [^1] More precisely Taillardat et al. (2023) construct their non-tail-equivalent random variable as $X:=Y\cdot\mathbb{1}(Y\leq u)+\left(Z+u\right)\cdot\mathbb{1}(Y>u)$, where $Z$ is a random variable with potentially very different tail behavior from $Y$. However the authors give no practical bounds on $u$, which might therefore be very large. Thus although it is always possible to construct non-tail-equivalent random variable it is not clear whether in practical applications $X$ will be markedly different from $Y$..
Nonetheless we believe Taillardat et al. (2023); Brehmer and Strokorb (2019) highlight an interesting question about the ability of the weighted and especially threshold-weighted CRPS to distinguish between tail regimes. We believe that there is value in studying this question empirically, thus in the following we consider a number of synthetic examples estimating distributions with different tail indices from samples.
@@ 7.2 Synthetic experiments
$Z\sim 0.4\cdot N(\mu,1)+0.6\cdot GPD(0,1,\xi).$ (19)
Here the $\mu$-parameter controls the location of the distribution body, whilst $\xi$ controls the tail structure. We fix the true $\mu=2$ and $\xi=0.5$. In the following we will consider experiments estimating both parameters using CRPS and twCRPS from 1000 samples of the true model ("training samples"). We consider a grid of candidate values $\hat{\mu},\hat{\xi}$ and for each of these we sample 250 values from the model in equation 19 and compute the sample average CRPS or twCRPS between those samples from the "candidate model" and the 1000 samples from the true model (the "training samples").
First we fix $\mu=2$ and only estimate the $\xi$ parameter, by varying it between 0.25 and 0.75, shown in Figure 9. All three estimation methods, using CRPS, twCRPS with a threshold based on the 90th percentile and twCRPS with a threshold based on the 95th percentile of the training distribution, seem to be able to stably identify the true tail parameter $\xi=0.5$. Interestingly however the twCRPS with 95th percentile threshold, which following theorem 4.1 effectively censors out 95% of the data seems to be just as stable for estimating $\xi$ as the CRPS. The loss curve seems to have a similar curvature around the minimum and does not seem more "noisy".
(a) CRPS
(b) twCRPS (90th percentile)
(c) twCRPS (95th percentile)
@@
In Figure 10 we estimate both parameters $\mu$ and $\xi$ using the CRPS and twCRPS with a 90th percentile threshold. We plot the CRPS and twCRPS marginally, such that for each value of $\mu$ the CRPS/twCRPS for all candidate values of $\xi$ can be seen and vice-versa. We find that the CRPS is much sharper for the estimation of $\mu$ than it is for the estimation of $\xi$, where the true parameter can’t be identified. This effect seems reversed for the twCRPS which is very inconclusive for $\mu$ estimation, but relatively sharp for the estimation of $\xi$. The twCRPS seems to focus the optimization on the parameter of interest, which controls the tail structure of the synthetic dataset, whilst the CRPS is concentrated on deviations in the body of the distribution.
(a) CRPS for values of $\mu$
(b) CRPS vor values of $\xi$
(c) twCRPS for values of $\mu$
(d) twCRPS for values of $\xi$
Figure 10: CRPS (top, a, b) and twCRPS with 90th percentile threshold (bottom, c, d) for different values of $\mu$ (left, a, c) and $\xi$ (right, b, d). The scatterpoints for each value of $\mu$ and $\xi$ correspond to all values of the respective other parameter. The true $\mu$ and $\xi$ are indicated by the blue vertical lines.
The effect of the twCRPS of focusing inference on the tail structure of a distribution follows mathematically from lemma 4.1. By using the twCRPS with threshold $\tau$ one effectively evaluates the censored CDF using the CRPS, meaning that the body of the distribution only factors in via censoring (the probability mass at the threshold $\tau$ itself). This explains the sharp $\xi$ inference in Figure 10, whilst the $\mu$ inference is inconclusive. For different applications different thresholds can be chosen by users depending on the application [^2] Especially when evaluating the construction in Taillardat et al. (2023) of $X:=Y\cdot\mathbb{1}(Y\leq u)+(Z+u)$ using the twCRPS one can vary the threshold and set it to $\tau>u$, which effectively centers inference onto $Z$. The weighted scoring rule in Taillardat et al. (2023) is assumed to be constant, however depending on the situation different weighting functions can be used for tailoring inference.. This process of using a CRPS for a censored distribution has some similarities with the usage of censored likelihood in extreme value theory (Huser and Davison, 2014) and it will be interesting to explore the usage of (weighted) proper scores for extreme value inference further in future work (see e.g. Yuen and Stoev 2014 for CRPS based inference in max stable processes and de Fondeville and Davison 2018 for inference using the gradient scoring rule).

### v4

@@ 6 Forecast combinations
These combine training using a CRPS and twCRPS and correspond to a weighted CRPS with weighting function $w(x)=1+\gamma\cdot\mathbb{1}(x\geq\tau)$, originally proposed by Thorarinsdottir and Schuhen (2018) for the assessment of forecasts. Other, non-threshold-based, weighting functions might also be usable to balance body-tail forecast performance, however, indicator weighting has the advantage of closed-form solutions being available. In a second step we consider linear pooling between the CRPS-trained and twCRPS-trained model, to control twCRPS improvement and CRPS loss. We focus on the 90th percentile threshold for both truncated normal and truncated logistic distributions.
@@ 6.1 Weighted combinations
We consider training using weighted scores of the form given in equation 16. In Figure 8 we show the CRPS skill over the CRPS-trained model (blue) and twCRPS skill over the twCRPS-trained model (orange) for different values of $\gamma$. For $\gamma=0$ the model corresponds to the CRPS-trained model whilst for growing $\gamma$ the twCRPS has stronger influence.
Varying $\gamma$ seems to allow users to control the CRPS loss and twCRPS improvement, with as $\gamma$ grows the resulting model approaches the twCRPS improvement of the twCRPS-trained model and its CRPS loss. Interestingly however, even though the twCRPS for the model with $\gamma=20$ is nearly at the level of the twCRPS-trained model (at least for the truncated normal distribution), the CRPS skill loss is still much better with only -1.58% CRPS skill loss for the $\gamma=20$ model against -3.64% for the twCRPS-trained models in the truncated normal case and only -1.75% against -2.85% in the truncated logistic case (see section 5.1). For cluster 4, as well as for larger values of $\gamma$ cluster 3 (and 2 in the truncated logistic cases) weighted training can even improve upon the “pure" twCRPS-model. Weighted training essentially leads to a more robust model, avoiding some cases where CRPS performance falls off strongly through twCRPS training, whilst still capturing much of the twCRPS improvement. This can also be of importance when training with very high thresholds, for example non-location-specific thresholds. When there are little to no observations over the threshold, twCRPS-based training can lead to highly mis-specified EMOS models, whilst weighted training allows to strike a compromise between the twCRPS and CRPS objectives (not shown).
The parameter $\gamma$ presents a “hyperparameter" that can be chosen by users to tailor a model’s CRPS and twCRPS performance for specific applications. However, after training, changing the balance of twCRPS improvement and CRPS loss requires retraining of the post-processing model. Thus in the following we will consider direct combination of CRPS-trained and twCRPS-trained models.
Figure 8: Weighted training: twCRPS skill score over the twCRPS-trained model (orange) and CRPS skill score over the CRPS-trained model (blue) as a function of $\gamma$ for both truncated normal (left) and truncated logistic (right) predictive distributions. Each diamond represents one cluster, whilst the lines correspond to the mean score.
@@ 6.2 Linear pool
The $\gamma$ parameter seems to “interpolate" between the CRPS-trained and twCRPS-trained model. We now study combining these models directly using their predictive CDFs. More concretely we study the following convex combination of the predictive distribution functions:
$F_{\text{pool}}=\lambda F_{\text{CRPS}}+(1-\lambda)F_{\text{twCRPS}}$ (17)
@@
This is also known as linear pool (Gneiting and Ranjan, 2013; Baran and Lerch, 2018) and the canonical case $\lambda=1/2$ corresponds to simple CDF averaging. The linear pool as distributional mixture provides a more flexible predictive distribution. It has some drawbacks, for example Gneiting and Ranjan (2013) show that it necessarily increases dispersion. More advanced versions such as the spread-adjusted linear pool have been proposed, however, we defer their study for future work.
@@
We study the twCRPS skill over the twCRPS-trained model and the CRPS skill over the CRPS-trained model, whilst varying $\lambda$ in Figure 9. The $\lambda$ parameter allows to strike a balance between the CRPS (being best for the CRPS-trained model and worst for the twCRPS-trained model) and the twCRPS (vice-versa). The curves are somewhat similar to the ones obtained by weighted training. As one can see the twCRPS improvements and CRPS losses have however very different “steepness", meaning that for example a model with $\lambda=0.6$ still has most of the twCRPS improvements, but controls the majority of the CRPS loss.
@@
Figure 9: Linear pool: twCRPS skill score over the twCRPS-trained model (orange) and CRPS skill score over the CRPS-trained model (blue) as a function of $\lambda$ for both truncated normal (left) and truncated logistic (right) distributions. Each diamond represents one cluster, whilst the lines correspond to the mean score.
@@
We investigate the example case $\lambda=0.6$ for the truncated normal distribution. Figure 10 shows the twCRPS and CRPS skill (relative to the CRPS-trained model) for both the twCRPS-trained and the linear pool model. As one can see the linear pool model keeps much of the twCRPS improvement, but controls the CRPS loss much stronger, with the spread across clusters being much smaller. Especially, for the CRPS performance, this will in part be due to the increased dispersion when linearly pooling forecasts. Pooling CRPS-trained and twCRPS-trained models seems to be able to somewhat alleviate the CRPS-twCRPS trade-off.
@@
Figure 10: Linear pool: twCRPS skill score (orange) and CRPS skill score (blue), both relative to the CRPS-trained model, for the twCRPS and the linear pool model with $\lambda=0.6$. Truncated normal predictive distribution, 90th percentile threshold.
@@
For different applications different choices of $\lambda$ will be relevant as users might be able to justify different CRPS losses. It might be possible to choose the $\lambda$ parameter in a data driven way, possibly based on the ensemble mean or external covariates such as weather patterns (Allen et al., 2021a; Spuler et al., 2024a, 2025). Simple experiments parametrizing $\lambda$ in terms of the distance of the ensemble mean to the threshold (a possible measure of the probability of exceedance), however did not show substantial skill over the ad-hoc choice of for example $\lambda=0.6$.
@@
Compared to the weighted training the linear pool directly interpolates between the two models, which means that for low values of $\lambda$ it is able to capture more of the twCRPS improvement, but also has much stronger CRPS loss. The latter is maximal with -1.58% for $\gamma=20$ for the weighted training, whilst reaching -3.64% for $\lambda=0$ in the linear pool, truncated normal case. By choosing adequate values of $\lambda$ and $\gamma$ however it is possible to construct models with very similar twCRPS and CRPS performance using both the linear pool and weighted training. Which one is preferable for applications might therefore depend on individual forecaster’s considerations, most notably whether changing the twCRPS/CRPS balance might be required during application. This would need retraining in the case of weighted training, making the linear pool preferable. For choosing $\gamma$ or $\lambda$ in applications there might be requirements on CRPS/twCRPS performance necessitating specific choices. Otherwise, we recommend an ad-hoc choice of a value that meets the objectives.
@@ 7 Interpreting the training impact of the twCRPS
After experiments studying the training of post-processing models using the threshold-weighted CRPS we now consider a number of synthetic experiments to understand the training impact of the twCRPS better in an idealized system. We focus on inferring tail properties from simulated data and on extreme value distributions, which have been used for various post-processing applications (see for example Baran and Lerch, 2015; Thorarinsdottir et al., 2016), to highlight how using a twCRPS can focus parameter inference on parameters controlling the distribution tail.
@@ 7.1 Background: Tail structures and proper scoring rules
Taillardat et al. (2023) argue that the mean CRPS and mean weighted CRPS cannot distinguish between non-tail-equivalent forecasts. They show that for any random variable $Y$ and any $\nu>0$ it is always possible to construct a non-tail-equivalent random variable $X$ (meaning a random variable with different tail index and thus differing tail behavior/return levels for extremes), such that:
$|\mathbb{E}_{y\sim Y}\left[\text{wCRPS}(F_{y},y)\right]-\mathbb{E}_{y\sim Y}\left[\text{wCRPS}(F_{x},y)\right]|\leq\nu$ (18)
where wCRPS is the weighted CRPS with any weighting function $w$. Brehmer and Strokorb (2019) generalise this and show that this is not only a property of the weighted CRPS, but rather all proper scoring rules fail to distinguish between tail equivalent forecasts in the sense of Taillardat et al. (2023). As Brehmer and Strokorb (2019) conclude this casts doubt on the ability of proper scoring rules to distinguish between tail regimes. In practical applications, however, proper scores such as the log-score (negative log-likelihood) are often successfully used to fit distributions to extremes and to distinguish between tail structures. This is supported by the theory of M-estimation, of which scoring rule inference is a special case (Dawid, 2007), providing consistency guarantees on parameter estimates. Thus the authors of this work believe that the arguments presented in Taillardat et al. (2023) and Brehmer and Strokorb (2019) do not carry conclusive practical implications. This is especially as Taillardat et al. (2023) do not provide bounds on the constructed non-tail-equivalent random variables $X$ that show that these will influence statistical analysis in practice  [^1] More precisely Taillardat et al. (2023) construct their non-tail-equivalent random variable as $X:=Y\cdot\mathbb{1}(Y\leq u)+\left(Z+u\right)\cdot\mathbb{1}(Y>u)$, where $Z$ is a random variable with potentially very different tail behavior from $Y$. However, the authors give no ($\nu$-dependent) bounds on $u$. $u$ might therefore be very large and even unobservable in practice. Thus, although it is always possible to construct a non-tail-equivalent random variable it is not clear whether in practical applications $X$ will be markedly different from $Y$. Furthermore, the weighted CRPS in equation 18 is assumed to be fixed. When using a twCRPS, however, the threshold can be adapted to the situation and can be set to $\tau>u$, which would effectively center inference onto $Z$ following Lemma 4.1.. Nonetheless, we believe Taillardat et al. (2023); Brehmer and Strokorb (2019) highlight an interesting question about the ability of the weighted and especially threshold-weighted CRPS to distinguish between tail regimes. We believe that there is value in studying this question empirically, thus in the following we consider a number of synthetic examples estimating distributions with different tail indices from samples.
@@ 7.2 Synthetic experiments
$0.4\cdot N(\mu,1)+0.6\cdot GPD(\xi),$ (19)
where
$F_{GPD}(x|\xi)=\begin{cases}1-\left(1+\xi x\right)^{-1/\xi}&\text{for }\xi\neq 0,\\ 1-\exp\left(-x\right)&\text{for }\xi=0.\end{cases}$
In this mixture the $\mu$-parameter controls the location of the distribution body, whilst $\xi$ controls the tail structure. We fix the true $\mu=2$ and $\xi=0.5$. In the following, we will consider experiments estimating both parameters using CRPS and twCRPS from 1000 samples of the true model (“training samples"). We consider a grid of candidate values $\hat{\mu},\hat{\xi}$ and for each of these we sample 250 values from the model in equation 19 and compute the sample average CRPS or twCRPS between those samples from the “candidate model" and the 1000 samples from the true model (the “training samples").
First we fix $\mu=2$ and only estimate the $\xi$ parameter, by varying it between 0.25 and 0.75, shown in Figure 11. All three estimation methods, using CRPS, twCRPS with a threshold based on the 90th percentile and twCRPS with a threshold based on the 95th percentile of the training distribution, seem to be able to stably identify the true tail parameter $\xi=0.5$. Interestingly however the twCRPS with 90th or 95th percentile thresholds, which following lemma 4.1 effectively censor out 90% / 95% of the data seems to be just as stable for estimating $\xi$ as the CRPS. The loss curves do not seem substantially more “noisy" and exhibit similar curvature across the domain of candidate $\xi$ values.
@@
In Figure 12 we estimate both parameters $\mu$ and $\xi$ using the CRPS and twCRPS with a 90th percentile threshold. We plot the CRPS and twCRPS marginally, such that for each value of $\mu$ the CRPS/twCRPS for all candidate values of $\xi$ can be seen and vice-versa. We find that the CRPS is much sharper for the estimation of $\mu$ than it is for the estimation of $\xi$, where the true parameter can’t be identified. This effect seems reversed for the twCRPS which is very inconclusive for $\mu$ estimation, but relatively sharp for the estimation of $\xi$. The twCRPS seems to focus the optimization on the parameter of interest, which controls the tail structure of the synthetic dataset, whilst the CRPS is concentrated on deviations in the body of the distribution. The type of proper score used for optimization can thus have an influence on the type of parameter focused during optimization.
Figure 12: CRPS (top, a, b) and twCRPS with 90th percentile threshold (bottom, c, d) for different values of $\mu$ (left, a, c) and $\xi$ (right, b, d). The scatterpoints for each value of $\mu$ and $\xi$ correspond to all values of the respective other parameter, the lines connect the values of the other parameter being constant. The true $\mu$ and $\xi$ are indicated by the blue vertical lines. The same random seed is held constant for all samples for all candidate values of $\mu$ and $\xi$.
The effect of the twCRPS of focusing inference on the tail structure of a distribution follows mathematically from lemma 4.1. By using the twCRPS with threshold $\tau$ one effectively evaluates the censored CDF using the CRPS, meaning that the body of the distribution only factors in via censoring (the probability mass at the threshold $\tau$ itself). This explains the sharp $\xi$ inference in Figure 12, whilst the $\mu$ inference is inconclusive. This process of using a CRPS for a censored distribution has some similarities with the usage of censored likelihood in extreme value theory (Huser and Davison, 2014) and it will be interesting to explore the usage of (weighted) proper scores for extreme value inference further in future work (see e.g. Yuen and Stoev 2014 for CRPS based inference in max stable processes and de Fondeville and Davison 2018 for inference using the gradient scoring rule).
Finally, both the CRPS and twCRPS are proper scoring rules, which means that in expectation both should be minimal at the true model. In practical applications, however, the true model is rarely within the candidate model class, and when optimizing a score for model training one only computes the in-sample average score, not the expected one. Using a weighted scores thus allows to account for some model deficiencies, but also allows more efficient estimation by focusing parameter inference, even if the true model is within the model class, as Fig 12 shows.